FOURIER TRANSFORM EMISSION SPECTROSCOPY OF ScS

\underline{R. S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ont., Canada N2L 3G1; and T. C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

The emission spectra of the A^{2}Π − X^{2}Σ^{+} and B^{2}Σ^{+} − X^{2}Σ^{+} systems of ScS have been observed in the 10000 − 13500 cm^{−1} region using a Fourier transform spectrometer. A rotational analysis of the 0 − 1, 0 − 0 and 1 − 0 bands of the A^{2}Π_{3/2} − X^{2}Σ^{+} and 0 − 1, 0 − 0, 1 − 0, 2 − 0 and 3 − 1 bands of the B^{2}Σ^{+} − X^{2}Σ^{+} systems has been obtained and improved spectroscopic constants have been extracted for the electronic states of ScS.